40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jul 12 23:15:34 2012, yuta, Update, LSC, PRMI LSC is making PRM motion worse PRMBSPRMIonoff.pngITMXITMYPRMIonoff.pngOplevPRMIonoff.png
    Reply  Sun Jul 15 16:25:15 2012, rana, Update, LSC, PRMI LSC is making PRM motion worse 
Message ID: 6972     Entry time: Thu Jul 12 23:15:34 2012     Reply to this: 6976
Author: yuta 
Type: Update 
Category: LSC 
Subject: PRMI LSC is making PRM motion worse 

It looks like PRMI LSC is making PRM motion (and BS motion) at ~3Hz worse.
I concluded this from measuring feedback signal of suspension servo and LSC servo.

Mechanism:
 1. BS and PRM moves alot at ~3 Hz.
 2. LSC senses fake signal at ~3Hz from beam spot motion on PD
 3. LSC feedback this motion to position of PRM
 4. Suspension damping servo try to cancel this because ~3 Hz motion is not actual length signal

Calculation:
x:   Orignal longitudinal motion of PRM
n_L: Sensing noise in LSC (including ITM motion, fake ~3Hz motion)
n_S: Sensing noise in suspension damping (OSEM sesor noise, fake ~3Hz motion)
G_L: Openloop transfer function of PRCL LSC
G_S: Openloop transfer function of suspension damping (PRM SUSPOS)
H:   LSC sensor transferfunction (PDH signal on REFL_33_I)
F_S: Filter for suspension damping
A:   Actuator transfer function (PRM OSEM coils)

  Since G_L >> G_S and G_L >> 1 for below 100Hz (see elogs #6950 and #6967), feedback signal of LSC and suspensiton damping can be written as

f_L = x - A*F_S*n_S - (1+G_S)/H*n_L
f_S = 1/G_L*x - A*F_S*n_S - G_S/H*n_L 

  So, basically, LSC supresses PRM motion but puts n_L to PRM. Suspension servo try to surpress n_L, which was not there when LSC is off.

Measurement:
 1. Below left is uncalibrated spectra of

Red:  suspension damping feedback to PRM/BS when PRMI is locked
Blue: LSC feeed back to PRM/BS when PRMI is locked
Pink: suspension damping feedback to PRM/BS when PRMI is not locked

  As you can see, PRM suspension damping feed back increases at ~ 1.5-3 Hz because of LSC. This is the same for BS at ~1 Hz and ~3 Hz.

PRMBSPRMIonoff.png    ITMXITMYPRMIonoff.png

 2. Above right is same spectra for ITMX/ITMY. There's no change in suspension damping feedback. This means, radiation pressure coupling or something is not related in this issue. LSC servo is not engaged for ITMs.

 3. Below is oplev spectra for PRM/BS

Red:  Oplev pitch error signal of PRM/BS when PRMI is locked
Blue: Oplev yaw error signal of PRM/BS to PRM/BS when PRMI is locked
Pink:  Oplev pitch error signal of PRM/BS when PRMI is not locked
Cyan: Oplev yaw error signal of PRM/BS to PRM/BS when PRMI is not locked

  You can see the increase in pitch/yaw motion at ~ 1.5-3 Hz for PRM, and ~1Hz/~3Hz for BS. They are consistent with measurement of feedback spectra.

OplevPRMIonoff.png



By the way:

  I adjusted oplev servo gains for ITMX. They were crazy this evening. They now have UGF ~ 2.5 Hz.

C1:SUS-ITMX_OLPIT_GAIN = 1.0 (was 2.6)
C1:SUS-ITMX_OLYAW_GAIN = -0.5 (was -1.6)


Next questions:
  - Can we notch ~3 Hz feedback so that LSC doesn't feedback this motion?
  - Why ~3 Hz motion is high for BS/PRM? Too much load on BS chamber stack?
  - Can we reduce ~3 Hz motion?
  - If BS chamber stack is bad, PR3 might have ~3 Hz motion, too. Does this make PRMI beam spot motion crazy?
  - How about PR2?

ELOG V3.1.3-